# Spin precession effects in the phasing formula of eccentric compact binary inspirals till the second post-Newtonian order

### v1

@@ II.3 Spin precession effects in TaylorT2
$\phi_{\rm SS,circ}=$ $\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{75\delta\kappa_{A}}{2}+\frac{96545\nu}{1984}+\biggl(75\nu-\frac{75}{2}\biggr)\kappa_{S}+\frac{1415}{3968}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{75\delta\kappa_{A}}{2}-\frac{99375\nu}{1984}$
@@
$+\biggl(75\nu-\frac{75}{2}\biggr)\kappa_{S}+\frac{1415}{3968}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[(150\nu-75)\kappa_{A}+\frac{1415\delta}{1984}-75\delta\kappa_{S}\biggr]$
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl[\biggl(\frac{525\nu}{67456}-\frac{525}{134912}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}-\frac{525\delta}{67456}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{525}{134912}$
$-\frac{525\nu}{67456}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}+\frac{525\delta}{67456}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\biggl(\frac{525\nu}{67456}-\frac{525}{134912}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+\biggl(\frac{525}{134912}$
@@
$-\frac{525\nu}{67456}\biggr)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{25\delta\kappa_{A}}{2}+\frac{2655\nu}{1984}+\biggl(\frac{25}{2}-25\nu\biggr)\kappa_{S}-\frac{2655}{3968}\biggr]$
@@
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[(25-50\nu)\kappa_{A}-\frac{2655\delta}{1984}+25\delta\kappa_{S}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{25\delta\kappa_{A}}{2}+\frac{2655\nu}{1984}+\biggl(\frac{25}{2}-25\nu\biggr)\kappa_{S}-\frac{2655}{3968}\biggr]\,$, (19e)
@@
$\phi_{\rm SS,ecc}=$ $\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{9345\delta\kappa_{A}}{8704}+\frac{117005\nu}{82688}+\biggl(\frac{9345\nu}{4352}-\frac{9345}{8704}\biggr)\kappa_{S}-\frac{455}{10336}\biggr]$
@@
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{9345\nu}{2176}-\frac{9345}{4352}\biggr)\kappa_{A}-\frac{455\delta}{5168}-\frac{9345\delta\kappa_{S}}{4352}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{9345\delta\kappa_{A}}{8704}-\frac{102445\nu}{82688}$
@@
$+\biggl(\frac{9345\nu}{4352}-\frac{9345}{8704}\biggr)\kappa_{S}-\frac{455}{10336}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{3115\delta\kappa_{A}}{8704}+\frac{1365\nu}{82688}+\biggl(\frac{3115}{8704}-\frac{3115\nu}{4352}\biggr)\kappa_{S}-\frac{1365}{165376}\biggr]$
@@
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{3115}{4352}-\frac{3115\nu}{2176}\biggr)\kappa_{A}-\frac{1365\delta}{82688}+\frac{3115\delta\kappa_{S}}{4352}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{3115\delta\kappa_{A}}{8704}+\frac{1365\nu}{82688}+\biggl(\frac{3115}{8704}$
@@
$-\frac{3115\nu}{4352}\biggr)\kappa_{S}-\frac{1365}{165376}\biggr]\biggr\}+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{18015\delta\kappa_{A}}{5632}+\frac{1435105\nu}{349184}+\biggl(\frac{18015\nu}{2816}-\frac{18015}{5632}\biggr)\kappa_{S}$
@@
$+\frac{73355}{698368}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{18015\nu}{1408}-\frac{18015}{2816}\biggr)\kappa_{A}+\frac{73355\delta}{349184}-\frac{18015\delta\kappa_{S}}{2816}\biggr]$
@@
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{18015\delta\kappa_{A}}{5632}-\frac{1581815\nu}{349184}+\biggl(\frac{18015\nu}{2816}-\frac{18015}{5632}\biggr)\kappa_{S}+\frac{73355}{698368}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{6005\delta\kappa_{A}}{5632}$
@@
$+\frac{54135\nu}{349184}+\biggl(\frac{6005}{5632}-\frac{6005\nu}{2816}\biggr)\kappa_{S}-\frac{54135}{698368}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{6005}{2816}-\frac{6005\nu}{1408}\biggr)\kappa_{A}-\frac{54135\delta}{349184}+\frac{6005\delta\kappa_{S}}{2816}\biggr]$
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{6005\delta\kappa_{A}}{5632}+\frac{54135\nu}{349184}+\biggl(\frac{6005}{5632}-\frac{6005\nu}{2816}\biggr)\kappa_{S}-\frac{54135}{698368}\biggr]\biggr\}+\biggl(\frac{y_{0}}{y}\biggr)^{50/3}\biggl\{\biggl(\frac{215775}{31966208}$
$-\frac{215775\nu}{15983104}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}+\frac{215775\delta}{15983104}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{215775\nu}{15983104}-\frac{215775}{31966208}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}$
$-\frac{215775\delta}{15983104}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\biggl(\frac{215775}{31966208}-\frac{215775\nu}{15983104}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+\biggl(\frac{215775\nu}{15983104}-\frac{215775}{31966208}\biggr)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr\}$ (19f)
@@ II.4 Resummation of the TaylorT2 phase
Note that in Eq. (20), the exponent of the $(1-e_{0}^{2})$ segment was arrived at by experimenting with various powers, and we finally settled with the one-fiftieth power since it proved to accumulate the smallest error in the number of cycles vis-a-vis the numerical solution.
@@
$\phi^{\rm(resum)}_{\rm SS,circ}$= $\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{75\delta\kappa_{A}}{2}+\frac{96545\nu}{1984}+\biggl(75\nu-\frac{75}{2}\biggr)\kappa_{S}+\frac{1415}{3968}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{75\delta\kappa_{A}}{2}-\frac{99375\nu}{1984}$
$+\biggl(75\nu-\frac{75}{2}\biggr)\kappa_{S}+\frac{1415}{3968}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[(150\nu-75)\kappa_{A}+\frac{1415\delta}{1984}-75\delta\kappa_{S}\biggr]$
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl[\biggl(\frac{525\nu}{67456}-\frac{525}{134912}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}-\frac{525\delta}{67456}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{525}{134912}$
$-\frac{525\nu}{67456}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}+\frac{525\delta}{67456}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\biggl(\frac{525\nu}{67456}-\frac{525}{134912}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+\biggl(\frac{525}{134912}$
$-\frac{525\nu}{67456}\biggr)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{25\delta\kappa_{A}}{2}+\frac{2655\nu}{1984}+\biggl(\frac{25}{2}-25\nu\biggr)\kappa_{S}-\frac{2655}{3968}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[(25-50\nu)\kappa_{A}-\frac{2655\delta}{1984}+25\delta\kappa_{S}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{25\delta\kappa_{A}}{2}+\frac{2655\nu}{1984}+\biggl(\frac{25}{2}-25\nu\biggr)\kappa_{S}-\frac{2655}{3968}\biggr]\,$, (22e)
$\phi^{\rm(resum)}_{\rm SS,ecc}$= $\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{3\delta\kappa_{A}}{4}+\frac{19309\nu}{19840}+\biggl(\frac{3\nu}{2}-\frac{3}{4}\biggr)\kappa_{S}+\frac{283}{39680}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{3\delta\kappa_{A}}{4}-\frac{3975\nu}{3968}$
@@
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{9345\delta\kappa_{A}}{8704}+\frac{1813777\nu}{1281664}+\biggl(\frac{9345\nu}{4352}-\frac{9345}{8704}\biggr)\kappa_{S}-\frac{226079}{5126656}\biggr]$
@@
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{9345\nu}{2176}-\frac{9345}{4352}\biggr)\kappa_{A}-\frac{226079\delta}{2563328}-\frac{9345\delta\kappa_{S}}{4352}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{9345\delta\kappa_{A}}{8704}$

### v2

@@ II.3 Spin precession effects in TaylorT2
$\phi_{\rm SS,circ}$ $=\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{75\delta\kappa_{A}}{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{22145\nu}{992}}+\biggl(75\nu-\frac{75}{2}\biggr)\kappa_{S}+\frac{1415}{3968}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{75\delta\kappa_{A}}{2}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{95\nu}{4}}+\biggl(75\nu-\frac{75}{2}\biggr)\kappa_{S}+\frac{1415}{3968}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[(150\nu-75)\kappa_{A}+\frac{1415\delta}{1984}-75\delta\,\kappa_{S}\biggr]$
@@
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{25\delta\kappa_{A}}{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{27455\nu}{992}}+\biggl(\frac{25}{2}-25\nu\biggr)\kappa_{S}-\frac{2655}{3968}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[(25-50\nu)\kappa_{A}-\frac{2655\delta}{1984}+25\delta\,\kappa_{S}\biggr]$
@@
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{25\delta\kappa_{A}}{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-25\nu}+\biggl(\frac{25}{2}-25\nu\biggr)\kappa_{S}-\frac{2655}{3968}\biggr]$
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl[\biggl({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{525\nu}{33728}}-\frac{525}{134912}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\left.-\frac{525\delta}{67456}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\right$.
$\left.{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{525}{134912}}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\right.\left.+\biggl(\frac{525}{134912}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{525\nu}{33728}}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\right$.
$\left.+\frac{525\delta}{67456}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\right.\left.{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{525}{134912}}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\right]\,$, (20a)
@@
$\phi_{\rm SS,ecc}$ $=\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{9345\delta\kappa_{A}}{8704}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{56455\nu}{82688}}+\biggl(\frac{9345\nu}{4352}-\frac{9345}{8704}\biggr)\kappa_{S}-\frac{455}{10336}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{9345\nu}{2176}-\frac{9345}{4352}\biggr)\kappa_{A}-\frac{455\delta}{5168}-\frac{9345\delta\kappa_{S}}{4352}\biggr]$
@@
$+\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{9345\delta\kappa_{A}}{8704}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2205\nu}{4352}}+\biggl(\frac{9345\nu}{4352}-\frac{9345}{8704}\biggr)\kappa_{S}-\frac{455}{10336}\biggr]$
@@
$+\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{A,0}\biggl[\frac{3115\delta\kappa_{A}}{8704}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{61915\nu}{82688}}+\biggl(\frac{3115}{8704}-\frac{3115\nu}{4352}\biggr)\kappa_{S}-\frac{1365}{165376}\biggr]$
$+\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{S,0}\biggl[\biggl(\frac{3115}{4352}-\frac{3115\nu}{2176}\biggr)\kappa_{A}-\frac{1365\delta}{82688}+\frac{3115\delta\kappa_{S}}{4352}\biggr]$
@@
$+\bm{\chi}_{S,0}\!\cdot\!\bm{\chi}_{S,0}\biggl[\frac{3115\delta\kappa_{A}}{8704}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{3115\nu}{4352}}+\biggl(\frac{3115}{8704}-\frac{3115\nu}{4352}\biggr)\kappa_{S}-\frac{1365}{165376}\biggr]\biggr\}$
@@
$+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{18015\delta\kappa_{A}}{5632}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{28925\nu}{15872}}+\biggl(\frac{18015\nu}{2816}-\frac{18015}{5632}\biggr)\kappa_{S}+\frac{73355}{698368}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{18015\nu}{1408}-\frac{18015}{2816}\biggr)\kappa_{A}+\frac{73355\delta}{349184}-\frac{18015\delta\kappa_{S}}{2816}\biggr]$
@@
$+\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{18015\delta\kappa_{A}}{5632}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{6315\nu}{2816}}+\biggl(\frac{18015\nu}{2816}-\frac{18015}{5632}\biggr)\kappa_{S}+\frac{73355}{698368}\biggr]$
@@
$+\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{A,0}\biggl[\frac{6005\delta\kappa_{A}}{5632}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{426445\nu}{174592}}+\biggl(\frac{6005}{5632}-\frac{6005\nu}{2816}\biggr)\kappa_{S}-\frac{54135}{698368}\biggr]$
$+\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{S,0}\biggl[\biggl(\frac{6005}{2816}-\frac{6005\nu}{1408}\biggr)\kappa_{A}-\frac{54135\delta}{349184}+\frac{6005\delta\kappa_{S}}{2816}\biggr]$
@@
$+\bm{\chi}_{S,0}\!\cdot\!\bm{\chi}_{S,0}\biggl[\frac{6005\delta\kappa_{A}}{5632}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{6005\nu}{2816}}+\biggl(\frac{6005}{5632}-\frac{6005\nu}{2816}\biggr)\kappa_{S}-\frac{54135}{698368}\biggr]\biggr\}$
$+\biggl(\frac{y_{0}}{y}\biggr)^{50/3}\biggl\{\biggl(\frac{215775}{31966208}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{215775\nu}{7991552}}\biggr)\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)^{2}+\frac{215775\delta}{15983104}\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)$
$+\frac{215775}{31966208}\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)^{2}+\biggl(-\frac{215775}{31966208}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{215775\nu}{7991552}}\biggr)\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{A,0}$
$-\frac{215775\delta}{15983104}\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{S,0}-\frac{215775}{31966208}\bm{\chi}_{S,0}\!\cdot\!\bm{\chi}_{S,0}\biggr\}.$ (20b)
Note that these coefficients up to $\mathcal{O}(e_{0}^{8})$ are given in [104].
@@ II.4 Resummation of the TaylorT2 phase
Note that in Eq. (21), the exponent of the $(1-e_{0}^{2})$ segment was arrived at by experimenting with various powers, and we finally settled with the one-fiftieth power since it proved to accumulate the smallest error in the number of cycles vis-a-vis the numerical solution. There is no particular physical justification for this choice.
@@
$\phi^{\rm(resum)}_{\rm SS,circ}$= $\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{75\delta\kappa_{A}}{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{22145\nu}{992}}+\biggl(75\nu-\frac{75}{2}\biggr)\kappa_{S}+\frac{1415}{3968}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{75\delta\kappa_{A}}{2}$-
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{95\nu}{4}}+\biggl(75\nu-\frac{75}{2}\biggr)\kappa_{S}+\frac{1415}{3968}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[(150\nu-75)\kappa_{A}+\frac{1415\delta}{1984}-75\delta\kappa_{S}\biggr]$
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl[\biggl({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{525\nu}{33728}}-\frac{525}{134912}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}-\frac{525\delta}{67456}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{525}{134912}$
$-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{525\nu}{33728}}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}+\frac{525\delta}{67456}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{525}{134912}}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+$
$-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{525}{134912}}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{25\delta\kappa_{A}}{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{27455\nu}{992}}+\biggl(\frac{25}{2}-25\nu\biggr)\kappa_{S}-\frac{2655}{3968}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[(25-50\nu)\kappa_{A}-\frac{2655\delta}{1984}+25\delta\kappa_{S}\biggr]$
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{25\delta\kappa_{A}}{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-25\nu}+\biggl(\frac{25}{2}-25\nu\biggr)\kappa_{S}-\frac{2655}{3968}\biggr]\,$, (23e)
$\phi^{\rm(resum)}_{\rm SS,ecc}$= $\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{3\delta\kappa_{A}}{4}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{4429\nu}{9920}}+\biggl(\frac{3\nu}{2}-\frac{3}{4}\biggr)\kappa_{S}+\frac{283}{39680}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{3\delta\kappa_{A}}{4}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{19\nu}{40}}$
@@
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{9345\delta\kappa_{A}}{8704}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1750903\nu}{2563328}}+\biggl(\frac{9345\nu}{4352}-\frac{9345}{8704}\biggr)\kappa_{S}-\frac{226079}{5126656}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{9345\nu}{2176}-\frac{9345}{4352}\biggr)\kappa_{A}-\frac{226079\delta}{2563328}-\frac{9345\delta\kappa_{S}}{4352}\biggr]$
@@
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{9345\delta\kappa_{A}}{8704}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2205\nu}{4352}}+\biggl(\frac{9345\nu}{4352}-\frac{9345}{8704}\biggr)\kappa_{S}-\frac{226079}{5126656}\biggr]$